2) A sample of 100 people is involved in a study for colorblindness. The results are summarized in the table below. An individual is selected at random from the 100 people. Let C the event that the person is colorblin Let F = the event that the person is female. Let M = the event that the person is male. Female Male Totals Colorblind 18 Not colorblind 54 22 60 40 24 76 100 Totals a.) P(Cn F)- b.) P(F) P(M I C)
Answer #1

a) from abvoe information:

P(C n F) =6/100 =0.06

b)P(F)=60/100 =0.6

c)P(M|C)=18/24 =3/4 =0.75

( please revert for any clarification or if there are other parts)